a capacitance of 20uf and a resistance of 10052 are connected in series across a 120v, 60Hz, if the capacitive reactance is 13352. find the impedance of the circuit. b). what is the magnitude of the circuit current? c). calculate the potential difference across the capacitor.


1
Expert's answer
2021-06-20T18:00:10-0400

Gives

C=20"\mu F"

R=10055 ohm

n=60Hz

V=120V

"X_c=13352\Omega"

Impedance of circuit


"Z=\sqrt{R^2+X^2_c}=\sqrt{10055^2+13352^2}=16714.63\Omega"

Current


"I=\frac{V}{Z}=\frac{120}{16714.63}=7.17\times10^{-3}A=7.17mA"

Potential difference across the capacitor

"V=I\times X_C"


"V_c=7.17\times10^{-3}\times13352 V=95.73\Omega"